OpenStudy (anonymous):

Help me solve this inequality please? will medal :3

OpenStudy (anonymous):

\[\left| x-1 \right|>7\]

OpenStudy (anonymous):

I got x>6, is that right?

OpenStudy (anonymous):

@danJS

pooja195 (pooja195):

You should have 2 answers

pooja195 (pooja195):

We know either x−1 > 7 or x−1 < −7

pooja195 (pooja195):

x−1>7(Possibility 1) x−1<−7(Possibility 2) Solve both of those for x

OpenStudy (anonymous):

i got 6 and -8

pooja195 (pooja195):

Can you show me how because the numbers are right but the symbols are wrong

OpenStudy (anonymous):

x-1>7 x-1>-7 -1 -1 -1 -1 x>6 x>-8

pooja195 (pooja195):

think of the two statements as x-1=7 x-1=-7 now solve for x

pooja195 (pooja195):

For both equations u would ADD the one to both sides rember we are isolating x

pooja195 (pooja195):

x-1>7 x-1>-7 +1 +1 +1 +1 x>? x>?

OpenStudy (anonymous):

ohhh i see what i did wrong v.v

pooja195 (pooja195):

thats good :) we are here to learn! :) now tell me what the answers are :)

OpenStudy (anonymous):

x>8 and x>-6

pooja195 (pooja195):

good :) x>8 or x<−6